Write the values of all four quantum numbers for 2p electrons. Discuss their
significance in brief.
1
Expert's answer
2019-02-07T03:11:50-0500

Four quantum numbers are used to describe the electron state in an atom: n, l, ml, ms

The principal quantum number (n) describes the electron shell, or energy level, of an electron. It describes the size and energy of an orbital.

The angular quantum number (l) dascribes the subshell. It describes the shape of the orbital. If l=0 the shepe of the orbital is spherical (s-orbital), if l=1 the shape of the orbital is polar (p-orbital), if l=2 the shape of the orbital is cloverleaf (d-orbital), if l=2 the shape of the orbital is more compicated (f-orbital). The angular quantum number (l) can be only integer btween o and n-1.

Magnetic quantum number (ml) describes the specific orbital within that subshell. It describes the orientation in space of a particular orbital. The magnetic quantim number can be any integer between -l and +l.

Spin quantum number describes the spin of the electron within an orbital. An electron has spin number s = 1/2, consequently ms = +1/2 and -1/2, referring to "spin up" and "spin down".

2p electrons are described with four quantum numbers:

n=2 (2nd electron shell)

l=n-1 = 2-1=1 (p orbital subshell)

ml = -1, 0, +1

ms = +1/2 and -1/2

As there are maximum 6 electrons on 2 p orbitals there are 6 sets of four quantum numbers:

1. n=2, l=1, ml = 1, ms = +1/2

2. n=2, l=1, ml = 1, ms = -1/2

3. n=2, l=1, ml = 0, ms = +1/2

4. n=2, l=1, ml = 0, ms = -1/2

5. n=2, l=1, ml = -1, ms = +1/2

6. n=2, l=1, ml = -1, ms = -1/2